# Spatiotemporal Change-Points in Development Discourse: Insights from Social Media in Low-Resource Contexts

### v1

@@ 1. Introduction
Social media platforms like X (formerly known as Twitter) have become central arenas for discourse about development, particularly in low-income countries where traditional media infrastructures and high-frequency, spatially granular development indicators remain limited (Gao et al., 2019; Sachs, 2006; Sen, 2014; Ferreira, 2024). These platforms function simultaneously as public spheres for debate and as living archives of collective sentiment, critique, and engagement with infrastructural change, policy shifts, and socio-political events (Karusala et al., 2019; Ngidi et al., 2016; Bergermann and Wolter, 2023), as well as sites for constructing and contesting development narratives (Dash et al., 2022). However, much of this scholarship has focused on short-lived spikes around elections, protests, or disasters, leaving the long-term spatial and temporal evolution of development-related discourse, especially in the Global South, underexplored.
@@
We contribute three advances to the scholarship of Human-Computer Interaction (HCI) and Information and Communications Technology for Development (ICTD). First, we conceptualize durable discourse as a new lens for understanding the endurance or decline of development concerns, that is, the extent to which issues persist or fade over time. This lens enables us to distinguish development issues that remain salient from those that diminish and strengthen citizen-driven sensing of public needs (Elwood, 2008; Barreneche and Lombana-Bermudez, 2023). Such continuity stands in contrast to the static nature of survey instruments. Second, we demonstrate that X functions as a sensitive detector of shifts in development agendas, providing empirical evidence that digital conversations both anticipate and accompany development flashpoints. Third, we introduce a framework that integrates mixed methods topic modeling and change-point detection, showing that geotagged social media data can yield contextualized insights in low-resource settings. By aligning digital traces with physical and policy interventions, this work demonstrates how temporally grounded and geospatially aware analysis can inform more responsive, equitable, and participatory development planning in Zambia and similar contexts.
@@ 2. Related Work
Development scholarship has evolved from narrow economic growth frameworks (Sachs, 2006) toward more holistic conceptualizations. Sen’s capabilities approach (Sen, 2014) reframes development as freedom, encompassing human capabilities, social transformation, and political empowerment as interconnected dimensions. This broader perspective views poverty not merely as income deficiency but as fundamental deprivation of well-being that constrains human flourishing. Contemporary development theory emphasizes the role of inclusive institutions in enabling broad-based economic and political participation (Robinson and Acemoglu, 2012; Acemoglu et al., 2015; Besley and Persson, 2010). This institutional perspective has driven a paradigm shift toward capacity-building and evidence-based policymaking, exemplified by innovations in rapid data collection for poverty measurement (Yoshida et al., 2020).
@@
2.1. Social Media and Development Discourse
@@
Emerging scholarship examines the digital engagement of marginalized populations with development themes. Karusala et al. (Karusala et al., 2019) explored regional disparities in autism awareness discourse between India and the United States through X hashtags, highlighting that information flows differ across development contexts. The #FeesMustFall movement in South Africa demonstrated social media’s role in anti-poverty activism and political mobilization (Ngidi et al., 2016), while analysis of global X discourse on the Rana Plaza collapse identified patterns of commemorative activity and international accountability demands (Bergermann and Wolter, 2023).
In India, X functions as an important arena for political participation (Jaidka and Ahmed, 2015) and a valuable lens for examining how politicians prioritize development concerns (Bozarth and Pal, 2019). Beyond its political and economic roles, the platform has also been employed to investigate social, human, and environmental development, including studies of COVID-19, violence, and natural disasters (Osuagwu et al., 2021; Jongman et al., 2015; Simon et al., 2014).
@@
These studies illuminate social media’s dual function as both a site of discourse and political action (c.f., (Dash et al., 2022)) in develping countries, underscoring the ways in which development narratives are constructed, contested, and circulated through digital platforms. Our work extends this literature by examining how such discourses evolve over extended timeframes in relation to physical infrastructure interventions.
@@ 2.2. Temporal Patterns in Social Media
In the development context specifically, Dhiman and Toshniwal (Dhiman and Toshniwal, 2022) demonstrated that X data can assess government campaign effectiveness in India, achieving strong correlations with official survey data while revealing spatial heterogeneity in public discourse across cities—an approach that aligns with our goal of connecting social media patterns to development outcomes at subnational scales.
2.3. Platform Equity and Representation
Critical platform studies scholarship has increasingly examined how dominant social media platforms marginalize underrepresented voices. Musgrave et al. (Musgrave et al., 2022) analyzed how Black women and femmes navigate harm and healing on social media, while Zhang et al. (Zhang et al., 2024) explored variations in platform trust across income groups. Similarly, social media users in the Global South are disproportionately concentrated in large urban centers and are more likely to be male and belong to specific language groups (Lasri et al., 2023; Mislove et al., 2011; Jiang et al., 2018), resulting in the oversampling of relatively advantaged populations (Hargittai, 2020).
Beyond open platforms like X, alternative social media ecosystems, particularly end-to-end encrypted messaging applications, have become dominant communication channels across the Global South. In Africa, WhatsApp has emerged as the default mode of telecommunication, favored both by government operatives seeking to spread messages surreptitiously and by opposition activists evading state controls (Cheeseman et al., 2020). Studies across Nigeria, Ghana, Kenya, Malawi, and Sierra Leone demonstrate that WhatsApp functions as a disruptive technology that simultaneously enables and undermines democratic consolidation, creating new opportunities for women and young people to engage in politics while also facilitating the spread of misinformation that exacerbates intergroup tensions (Cheeseman et al., 2020).
@@ 3.3. Qualitative Analysis
Sen’s capability approach defines poverty as a deprivation of capabilities, not merely a lack of income. Multidimensional approaches build on this by offering a broader perspective beyond monetary metrics (Alkire and Santos, 2010). Aligned with these concepts, the Sustainable Development Goals (SDGs) provide a global framework for interconnected social, economic, and environmental challenges. By interpreting the final topics through the lenses of multidimensional poverty and the SDGs, we can better articulate how the X discourse in Zambia reflects the overlapping issues of poverty and regional development.
@@ 3.5.1. Choice of Region and Event
We chose to focus on two regions that were critically affected by different events unfolding at different paces, one acute (COVID in the most populace province, Lusaka), the other deliberate and planned (infrastructure project in Southern province).
@@ 4.1. Development Topics
Our analysis surfaced seven development-related theme: election corruption, food systems, social progress, mining, frustration with government policy, public health challenges, and social inequality. These seven themes map closely onto core Sustainable Development Goals (SDGs): election corruption and frustration with government policy relate to SDG 16 (Peace, Justice, and Strong Institutions); food systems aligns with SDG 2 (Zero Hunger); social progress with SDGs 5 and 10 (Gender Equality; Reduced Inequalities); mining with SDG 8 (Decent Work and Economic Growth); public health challenges with SDG 3 (Good Health and Well-Being); and social inequality with SDGs 1 and 10 (No Poverty; Reduced Inequalities). Framing the themes in relation to the SDGs emphasizes that poverty, as expressed in the sampled tweets, extends far beyond income to encompass political, social, and environmental dimensions.
@@ 4.2. Change-Points Over Time
This shift exemplifies how acute public health emergencies can eclipse sustained attention to endemic diseases and preventive health infrastructure. Zambia has long grappled with cholera outbreaks, particularly in urban Lusaka (Gulumbe et al., 2024), yet pandemic discourse temporarily displaced these ongoing concerns. The significant downward trend after June 14 (p = 0.02) suggests that COVID-related attention, while intense, was relatively short-lived—a pattern consistent with crisis communication research showing rapid attention decay following initial emergency phases.
@@
Unlike the acute, short-lived public health spike observed in Lusaka, this policy discourse shift represents a durable reorientation. The PELT model identifies an extended post-change regime stretching well beyond the initial March 29 break, indicating that the geothermal project catalyzed a lasting transformation in how Southern Province residents discussed government policy—moving from tourism-centric concerns to sustained engagement with infrastructure governance and resource management.
@@ 5. Discussion
Given the limited ground-truth data associated with wealth in developing countries in which traditional poverty assessments focus on aggregated data at lower-spatial resolutions (Jung, 2023) and missing data in developing country census-based measurements (Carr-Hill, 2013; Kuffer et al., 2022), our work illustrates how local change-points in time series showing flashpoints of interest associated with policy challenges (e.g., COVID-19) and new policy interventions (e.g., new infrastructure projects like the geothermal project discussed in Southern province). These spatio-temporal focal points can provide government planners with foresight (Kayser and Bierwisch, 2016) as they study the needs of citizens on the ground. A responsive government can sense salient issues raised by networked publics (Büscher et al., 2014) on social media to better understand the needs of the population and respond to them in a timely fashion.

### v2

@@ 1. Introduction
Social media platforms like X (formerly known as Twitter) have become central arenas for discourse about development, particularly in low-income countries where traditional media infrastructures and high-frequency, spatially granular development indicators remain limited (Gao et al., 2019; Sachs, 2006; Sen, 2014; Ferreira, 2024). These platforms function as designed spaces—what scholarship has termed “networked public spheres” (Semaan et al., 2014; Balestrini et al., 2017)—where citizens construct, contest, and sustain narratives about development conditions. Unlike traditional surveys that provide discrete snapshots, digital discourse offers continuous signals of public priorities. Yet while HCI research has extensively examined how to design for citizen engagement (Cazacu et al., 2021; Vlachokyriakos et al., 2016), less attention has been paid to measuring the temporal dynamics of such engagement: How long do citizen concerns persist? When does discourse crystallize into sustained public attention versus dissipating after initial interest?
@@
We contribute three advances to the scholarship of Human-Computer Interaction (HCI), Information and Communications Technology for Development (ICTD), and digital civics. First, we operationalize durable discourse as a measurable framework for understanding the persistence of citizen voice in digital public squares, drawing on crisis informatics’ temporal models (Reuter and Kaufhold, 2018; Soden and Palen, 2018) to distinguish development issues that sustain public attention from those that fade. Second, we demonstrate that X functions as a sensitive detector of shifts in development agendas, providing empirical evidence that digital conversations both anticipate and accompany development flashpoints. Third, we introduce design implications for development monitoring systems that align with DIS scholarship on civic technology design (Corbett and Le Dantec, 2019; Fox and Le Dantec, 2014), showing how temporally grounded analysis can inform more responsive and participatory development planning in Zambia and similar contexts.
1.1. Development as Citizen Agency
Sen’s capability approach reframes poverty not as income deficiency but as deprivation of well-being that constrains human flourishing (Sen, 2014; Jacobson, 2016). Central to this framework is the concept of citizen agency—the idea that citizens, not governments or external development agencies, should articulate which functionings they value and which capabilities they seek (Jacobson, 2016). Sen emphasizes that public speech plays a central role in the processes through which citizen agency drives collective identification of development priorities. This theoretical orientation positions social media discourse not merely as data to be extracted but as communicative agency in action—citizens voicing development concerns through digital means.
Multidimensional approaches build on this by offering a broader perspective beyond monetary metrics (Alkire and Santos, 2010). Aligned with these concepts, the Sustainable Development Goals (SDGs) provide a global framework for interconnected social, economic, and environmental challenges. By interpreting development discourse through the lenses of multidimensional poverty and the SDGs, we can better articulate how citizen voice on social media reflects the overlapping issues of poverty and regional development.
1.2. The Digital Public Square
Digital civics scholarship examines how technologies mediate relationships between citizens, communities, and governance structures (Vlachokyriakos et al., 2016; Cazacu et al., 2021). Balestrini et al. (Balestrini et al., 2017) demonstrated how designed systems can orchestrate citizen engagement around urban issues, while Semaan et al. (Semaan et al., 2014) showed that social media can support political deliberation across multiple public spheres. These platforms function as what DiSalvo et al. (DiSalvo et al., 2014) term sites for expressing “matters of concern”—issues that gather publics around shared problems.
Social media platforms thus serve dual functions: as archives of collective sentiment and as living sites where development narratives are constructed, contested, and circulated (Karusala et al., 2019; Ngidi et al., 2016; Bergermann and Wolter, 2023; Dash et al., 2022). However, much of this scholarship has focused on short-lived spikes around elections, protests, or disasters, leaving the long-term spatial and temporal evolution of development-related discourse, especially in the Global South, underexplored. Measuring the temporal dynamics of public formation remains challenging: When do citizen concerns achieve the “gathering” function that creates sustained public pressure? When does discourse remain ephemeral, dissipating after initial attention? These questions motivate our introduction of durable discourse as a measurement framework.
1.3. Operationalizing Durable Discourse
Crisis informatics scholarship has long recognized that crises produce both acute (immediate) and chronic (long-term) effects on communities, yet research has predominantly focused on the immediate response phase (Soden and Palen, 2018; Randazzo et al., 2025). Established temporal models in disaster management distinguish multiple phases: the acute crisis period (days to two weeks), emergency response and early recovery (days to three months), short-term recovery (three months to two years), and long-term reconstruction (two to ten years) (National Research Council, 2006; Reuter and Kaufhold, 2018). Similarly, psychological models identify a “honeymoon phase” of social cohesion immediately following disaster, followed by a “disillusionment phase” when communities confront long-term repercussions (De Silva et al., 2023). Yet measuring when discourse transitions between these phases—and which concerns persist versus fade—remains methodologically challenging.
Recent scholarship has advanced spatiotemporal methods for analyzing social media during crises. Yang et al. (Yang et al., 2016) developed the SMART testbed for exploring spatiotemporal patterns in geo-targeted social media, demonstrating how crisis management applications can track human dynamics across space and time. Similarly, other studies have examined how trauma discourses unfold on Twitter following crisis events, revealing the cultural negotiation processes through which communities make sense of disruption (Eriksson, 2018). Mirbabaie et al. (Mirbabaie et al., 2020) introduce the concept of “sense-giving” in crisis communication, showing how authoritative actors can maintain influence over extended crisis discourse—a temporal dynamic our durable discourse framework aims to capture and measure.
We define durable discourse as sustained public engagement with development issues that persists beyond initial event-driven attention spikes. This concept bridges crisis informatics’ temporal frameworks with design research’s interest in how technologies mediate civic engagement over time. We operationalize durability through three measurable criteria:
  1. (1)
    Temporal persistence: Discourse that maintains elevated engagement levels for 60+ days following a significant change-point, as detected by the PELT algorithm. This threshold aligns with crisis research distinguishing acute response (under 60 days) from sustained recovery engagement (Reuter and Kaufhold, 2018).
  2. (2)
    Thematic coherence: Sustained discourse that maintains semantic consistency rather than fragmenting into unrelated sub-topics, assessed through topic coherence scores and qualitative validation.
  3. (3)
    Structural signature: A change-point pattern characterized by (a) statistically significant shift (p ¡ 0.10 via Mann-Kendall test), (b) extended post-change regime versus rapid return to baseline, and (c) vocabulary transformation (via log-likelihood ratio analysis) indicating thematic reorientation rather than mere volume changes.
This operationalization enables us to distinguish ephemeral discourse—intense but short-lived attention spikes characteristic of acute crisis phases that demand immediate response—from durable discourse—persistent engagement patterns characteristic of structural concerns that warrant longer-term policy attention. As Soden and Palen (Soden and Palen, 2018) argue, crisis informatics must attend to the “long now” of disasters—the extended temporal horizons over which communities make sense of and recover from disruptive events. Our framework extends this insight to development contexts, providing measurement criteria for identifying when citizen concerns achieve sustained public attention. Table 1 summarizes these distinctions with examples from our analysis.
Table 1. Distinguishing Ephemeral from Durable Discourse
Characteristic Ephemeral Durable
Duration <<60 days elevated >>60 days elevated
Trajectory Sharp peak, rapid decline Sustained plateau
Policy response Immediate crisis management Long-term engagement
Example COVID-19 in Lusaka Geothermal Infrastructure Project in Southern
@@ 2. Related Work
Development scholarship has evolved from narrow economic growth frameworks (Sachs, 2006) toward more holistic conceptualizations. Contemporary development theory emphasizes the role of inclusive institutions in enabling broad-based economic and political participation (Robinson and Acemoglu, 2012; Acemoglu et al., 2015; Besley and Persson, 2010). This institutional perspective has driven a paradigm shift toward capacity-building and evidence-based policymaking, exemplified by innovations in rapid data collection for poverty measurement (Yoshida et al., 2020).
@@
2.1. Digital Civics and Measuring Community Discourse
Digital civics research within HCI examines how technologies enable citizen voice in civic processes (Vlachokyriakos et al., 2016; Cazacu et al., 2021). This scholarship emphasizes participatory approaches where communities actively shape governance outcomes through designed systems. At DIS specifically, Corbett and Le Dantec (Corbett and Le Dantec, 2019) developed frameworks for designing civic technology with trust, while Fox and Le Dantec (Fox and Le Dantec, 2014) examined how technologies can scaffold community engagement through culture and heritage. Crivellaro et al. (Crivellaro et al., 2014) examined how Facebook enabled political discourse and the emergence of social movements, while Dow et al. (Dow et al., 2018) explored tensions between grassroots participation and institutional hierarchies in public services.
Recent DIS work has focused on how platform features shape political discussions. Research on affirmative action discourse across Reddit, Twitter/X, and TikTok (Pyle et al., 2025) demonstrates how visibility and association affordances shape how users present evidence and identity cues in civic discourse—highlighting that platform design fundamentally shapes how citizen voice is expressed and received. Similarly, Schroeter et al. (Schroeter et al., 2012) explored how urban screens can be designed for situated public engagement, while Gordon and Manosevitch (Gordon and Manosevitch, 2011) demonstrated how augmented deliberation can merge physical and virtual interaction to engage communities in urban planning.
Nielsen et al. (Holten Holten Møller et al., 2021) examined how unemployed individuals navigate data sharing in bureaucratic decision-making—revealing how citizens exercise agency in data-driven public services. This work underscores that citizen voice in digital contexts is not merely about expression but about negotiating power relations with institutions. Our work contributes to this literature by introducing measurement approaches for the temporal persistence of citizen voice. Where prior work has examined how platforms enable civic engagement, we focus on measuring how such engagement unfolds and persists over time—providing a complementary lens for understanding when citizen concerns crystallize into sustained public attention.
2.2. Inequality and Discourse in Developing Digital Contexts
Critical platform studies scholarship has increasingly scrutinized the mechanisms through which dominant social media platforms marginalize underrepresented voices. For instance, Musgrave et al. (Musgrave et al., 2022) analyzed how Black women and femmes navigate harm and healing on social media, while Zhang et al. (Zhang et al., 2024) explored significant variations in platform trust across income groups. These inequalities are particularly pronounced in the Global South, where users are disproportionately concentrated in large urban centers and are more likely to be male and belong to specific language groups (Lasri et al., 2023; Mislove et al., 2011; Jiang et al., 2018). This demographic skew results in the oversampling of relatively advantaged populations on open platforms (Hargittai, 2020).
In response to these exclusionary dynamics, alternative social media ecosystems—particularly end-to-end encrypted messaging applications—have become dominant communication channels. In Africa, WhatsApp has emerged as the default mode of telecommunication, favored both by government operatives seeking to spread messages surreptitiously and by opposition activists evading state controls (Cheeseman et al., 2020). Studies across Nigeria, Ghana, Kenya, Malawi, and Sierra Leone demonstrate that WhatsApp functions as a disruptive technology that simultaneously enables and undermines democratic consolidation, creating new opportunities for women and young people to engage in politics while also facilitating the spread of misinformation that exacerbates intergroup tensions (Cheeseman et al., 2020).
@@ 2.2. Inequality and Discourse in Developing Digital Contexts
Parallel to these platform-specific investigations, emerging scholarship has begun to examine how marginalized populations utilize digital tools to engage specifically with development themes. Karusala et al. (Karusala et al., 2019) explored regional disparities in autism awareness discourse between India and the United States through X hashtags, highlighting that information flows differ markedly across development contexts. Similarly, the #FeesMustFall movement in South Africa demonstrated social media’s role in anti-poverty activism and political mobilization (Ngidi et al., 2016), while analysis of global X discourse on the Rana Plaza collapse identified patterns of commemorative activity and international accountability demands (Bergermann and Wolter, 2023).
@@
This intersection of political participation and development discourse is particularly visible in India, where X functions as both an important arena for political participation (Jaidka and Ahmed, 2015) and a valuable lens for examining how politicians prioritize development concerns (Bozarth and Pal, 2019). Beyond its political and economic roles, the platform has also been employed to investigate social, human, and environmental development, including studies of COVID-19, violence, and natural disasters (Osuagwu et al., 2021; Jongman et al., 2015; Simon et al., 2014). Collectively, these studies illuminate social media’s dual function as a site of both discourse and political action (c.f., (Dash et al., 2022)) in developing countries, underscoring the ways in which development narratives are constructed, contested, and circulated through digital platforms. Our work extends this literature by examining how such discourses evolve over extended timeframes in relation to physical infrastructure interventions.
@@ 3.5.1. Choice of Region and Event
We chose to focus on two regions that were critically affected by different events unfolding at different paces, one acute (COVID in the most populace province, Lusaka), the other deliberate and planned (infrastructure project in Southern province). These cases allow us to empirically test our operationalization of durable discourse (see §1.3) by comparing ephemeral crisis-driven attention with sustained structural engagement.
@@ 4.1. Development Topics
Our analysis surfaced seven development-related themes: election corruption, food systems, social progress, mining, frustration with government policy, public health challenges, and social inequality. These seven themes map closely onto core Sustainable Development Goals (SDGs): election corruption and frustration with government policy relate to SDG 16 (Peace, Justice, and Strong Institutions); food systems aligns with SDG 2 (Zero Hunger); social progress with SDGs 5 and 10 (Gender Equality; Reduced Inequalities); mining with SDG 8 (Decent Work and Economic Growth); public health challenges with SDG 3 (Good Health and Well-Being); and social inequality with SDGs 1 and 10 (No Poverty; Reduced Inequalities). Framing the themes in relation to the SDGs emphasizes that poverty, as expressed in the sampled tweets, extends far beyond income to encompass political, social, and environmental dimensions.
@@ 4.2. Change-Points Over Time
This shift exemplifies how acute public health emergencies can eclipse sustained attention to endemic diseases and preventive health infrastructure. Zambia has long grappled with cholera outbreaks, particularly in urban Lusaka (Gulumbe et al., 2024), yet pandemic discourse temporarily displaced these ongoing concerns. The significant downward trend after June 14 (p = 0.02) suggests that COVID-related attention, while intense, was relatively short-lived—a pattern consistent with crisis communication research showing rapid attention decay following initial emergency phases. Using our operationalization from §1.3, this constitutes ephemeral discourse: the attention spike lasted fewer than 60 days before declining, with rapid return toward baseline engagement levels.
@@
Unlike the acute, short-lived public health spike observed in Lusaka, this policy discourse shift represents a durable reorientation. The PELT model identifies an extended post-change regime stretching well beyond the initial March 29 break, indicating that the geothermal project catalyzed a lasting transformation in how Southern Province residents discussed government policy—moving from tourism-centric concerns to sustained engagement with infrastructure governance and resource management. Using our operationalization from §1.3, this constitutes durable discourse: the thematic reorientation persisted for more than 60 days with sustained elevated engagement and maintained semantic coherence around infrastructure and resource governance themes.
@@ 5. Discussion
Our analysis demonstrates how temporally-grounded discourse analysis can surface citizen voice in development contexts. By distinguishing ephemeral from durable discourse patterns using our operationalized criteria (§1.3), we provide a measurement framework that complements traditional development indicators. The COVID-19 case illustrates ephemeral discourse—intense but short-lived engagement characteristic of what crisis informatics terms the “acute phase” (Reuter and Kaufhold, 2018). The geothermal case illustrates durable discourse—sustained engagement with infrastructure governance that persisted well beyond the initial announcement, characteristic of what Soden and Palen (Soden and Palen, 2018) call the “long now” of disruptive events. This distinction carries significant implications for development practice: crisis-driven discourse requires rapid organizational response, while durable discourse warrants sustained stakeholder engagement strategies.
Given the limited ground-truth data associated with wealth in developing countries in which traditional poverty assessments focus on aggregated data at lower-spatial resolutions (Jung, 2023) and missing data in developing country census-based measurements (Carr-Hill, 2013; Kuffer et al., 2022), our work illustrates how local change-points in time series can reveal flashpoints of interest associated with policy challenges (e.g., COVID-19) and new policy interventions (e.g., new infrastructure projects like the geothermal project discussed in Southern province). These spatio-temporal focal points can provide government planners with foresight (Kayser and Bierwisch, 2016) as they study the needs of citizens on the ground. A responsive government can sense salient issues raised by networked publics (Büscher et al., 2014) on social media to better understand the needs of the population and respond to them in a timely fashion.
@@ 5.3. Implications for AI-Supported Development Practice
5.4. Design Implications for Development Monitoring Systems
Our findings suggest several implications for the design of systems that seek to surface citizen voice in development contexts. These implications align with DIS scholarship emphasizing that design must account for the systemic, temporal, and political dimensions of interactive systems (Corbett and Le Dantec, 2019; Fox and Le Dantec, 2014):
DI1: Temporal differentiation. Monitoring systems should distinguish between ephemeral and durable discourse patterns, as these require different organizational responses. Crisis-driven discourse (like our COVID-19 case) warrants rapid response protocols and immediate communication adjustments—aligning with crisis informatics’ acute phase interventions (Reuter and Kaufhold, 2018). Durable discourse (like infrastructure concerns) indicates sustained public concern requiring longer-term stakeholder engagement strategies characteristic of recovery and reconstruction phases. Systems should surface these temporal signatures to practitioners through differentiated visual encodings or alert mechanisms.
DI2: Regional granularity. Our change-point analysis revealed distinct temporal signatures across provinces—the same development theme (frustration with government policy) exhibited different dynamics in Lusaka versus Southern Province. Development monitoring systems should support subnational filtering to capture local discourse dynamics that aggregate national analyses would obscure. This aligns with DIS work on situated engagement (Schroeter et al., 2012), which emphasizes that civic technologies must account for place-based differences in how communities engage with issues.
DI3: Vocabulary tracking. The LLR analysis revealed that change-points involve not just volume shifts but vocabulary transformations. Systems should track semantic evolution—the shift from “cholera, elimination, systems” to “safe, mask, tests”—not just topic prevalence, to detect genuine thematic reorientations versus mere attention spikes. This design consideration supports what Corbett and Le Dantec (Corbett and Le Dantec, 2019) describe as designing for trust—ensuring that systems accurately represent citizen concerns rather than amplifying noise.
DI4: Multi-platform integration. Given the platform equity concerns we discuss (see §LABEL:sec:whatsapp), monitoring systems should be designed to integrate signals from multiple platforms rather than relying solely on any single source. In Global South contexts where WhatsApp dominates, designs should accommodate privacy-preserving data donation approaches alongside open platform analysis. This extends DIS scholarship on designing for marginalized communities (Fox and Le Dantec, 2014) by acknowledging that different populations express civic voice through different technological channels.
5.5. Practitioner Use Cases
To illustrate how our framework might inform practice, we describe three hypothetical use cases grounded in our findings:
Use Case 1: Ministry of Health Crisis Communication. When our PELT model detects a significant upward change-point in public health discourse (as in the Lusaka COVID-19 case), health ministry communication teams could use the associated LLR vocabulary analysis to tailor public messaging. The shift from “cholera, elimination, systems” to “safe, mask, tests” indicates public attention has pivoted to immediate protective measures—suggesting communication should emphasize actionable safety guidance rather than long-term health planning rhetoric.
Use Case 2: Infrastructure Planning Agency. The sustained discourse shift around the geothermal project in Southern Province—with vocabulary moving from tourism (“resorts, victoria”) to resource governance (“resources, decisions, legal”)—signals that communities are engaging with infrastructure development not just as economic opportunity but as a governance concern. Planners could use such signals to proactively address transparency and community consultation, anticipating that these concerns will persist rather than fade.
Use Case 3: NGO Program Targeting. Development NGOs could use topic-based wealth prediction (our $R^{2}=0.610$ finding in Appendix A) to complement traditional needs assessments. Where survey data is outdated or unavailable, discourse patterns could provide interim signals for resource allocation—with the important caveat that Twitter data amplifies already-connected populations, necessitating triangulation with other data sources (see §6 for more details).